A share of TRS, Inc. preferred stock has a Par Value of $100 and a Dividend Rate of 13%. If you desire a return on investment of 19%, what is the price you would pay for the stock?
Answer #1

Annual dividend = 0.13 * 100 = 13

Price of stock = Annual dividend / required return

Price of stock = 13 / 0.19

Price of stock = $68.42